given Y=C+I+G+X

C=100+0.9Yd I=200-500r M=0.8Y-2000r X=100-0.2Y-500r G=200 T=0.2 L=800

Compute the values of C,L,X and M

Calculate the monetary and fiscal policy multipliers and interpret them


1
Expert's answer
2021-08-06T15:27:27-0400

Consumption Function "C=100+0.9Yd"

Investment Function"I=200-500r"

Real Money supply "M=0.8Y-2000r"

Net Export"X=100-0.2Y-500r"

Government Spending G=200

Tax Rate T=0.2

Real Money demanded L=800


"Y=C+I+G+X\\

 = 100+0.9Yd+200-500r+200+100-0.2Y-500r\\

 = 600 + 0.9(0.8y) - 1000r -0.12y\\

 = 600 + 0.6Y -1000r"


"0.4Y = 600 - 1000r"

  "Y = 1500 - 2500r" ..................Equation for IS

At equilibrium Money supply is equal to the money demanded

"Ms = Md\\

Y - 100 r = 800"

"Y = 800 + 100r" .............................Equation for LM

 

We Now AT equilibrium IS curve Intersect LM curve so,

"1500 + 2500r = 800+ 100r\\

 r = \frac{7}{26}"

Thus,

"Y = 800 + 100(\frac{7}{26})\\

 = \frac{21500}{26}"


"C = 100 + 0.72Y\\

 = 100 + 0.72(\frac{21500}{26})\\

 =100 + 577\\

 = 677"


"I = 200 - 500(\frac{7}{26})\\

 = \frac{1700}{26}"


"X = 100 -0.12(\frac{21500}{26}) - 500(\frac{7}{26})\\

 =\frac{-3480}{26}"